Raymond is a basketball player who takes four independent free throws with 70% probability of getting a basket on each shot. Let X be the number of baskets Raymond gets. Find the probability that he gets exactly 2 baskets, to 3 decimal places.
1
Expert's answer
2020-06-02T20:01:45-0400

 Let the Probability of success be that the player get the basket in a shot which is given as 0.7 (70%).

X is the number of shots which is given that player takes four independent free throws.

Using, Binomials formula:

"P( x = r) = \begin{pmatrix}
 n \\
 r 
\end{pmatrix}* P^r( 1 - P)^{n - r}"

where P is the Probability of success, n is number of trials.

In our case , n = 4 ; r =2; P = 0.7

P( exactly 2 baskets) = P(x = 2) "= \begin{pmatrix}
 4 \\
 2
\end{pmatrix}*(0.7)^2* ( 1 - 0.7)^{4 - 2} ="

= "[ 4!\/ 2!(4-2)!] (0.7)^2 ( 1 - 0.7)^{4 - 2} ="

"= (4*3\/2)* (0.7)^2* ( 0.3)^2 ="

"= 6* ( 0.49)* (0.09)"


P( x= 2) = 0.265

Answer: Probability that Raymonds gets exactly 2 baskets is 0.265.